# Solving Truly Massive Budgeted Monotonic POMDPs with Oracle-Guided Meta-Reinforcement Learning

### v1

@@ 6 Conclusions
In this paper we introduce a framework to provide computationally efficient solutions, scalable for large scale budget-constrained multi-component monotonic POMDPs. Noting that the coupling is expressed solely through the shared budget, our method functions by first decomposing the large $n$-component POMDP into $n$ independent single-component POMDPs using a random forest budget allocation of the shared budget. Next, we compute the oracle policy using value iteration for each component POMDP and budget pair. Finally, using this oracle policy together with a meta-trained PPO agent, we solve each component POMDP. Consequently, we obtain the approximately optimal policy for the multi-component POMDP. The experimental evaluations conducted for a real-world scenario of large scale infrastructure maintenance show that the proposed approach outperforms existing baselines and almost matches the performance of the oracle policy, which relies on information not available to the planner in a partially observable setting. Furthermore, computational complexity analysis demonstrates the scalability of our framework. Future work will focus on extending the framework’s capabilities to more dynamic budget allocation schemes and more complicated hierarchical budget constraints.

### v2

@@ 6 Conclusions
We proposed a scalable framework for solving budget-constrained multi-component monotonic POMDPs. Our chief theoretical contribution is a proof that the single-component value function is concave in budget, which underpins an efficient two-step solution strategy. First, a random-forest surrogate exploits that concavity to distribute the shared budget across components, thereby decomposing the large $n$-component POMDP into $n$ independent single-component POMDPs. Second, an oracle-guided, meta-trained PPO agent—shaped by value iteration on the fully observable counterpart—learns a near-optimal policy for each component–budget pair. Comprehensive experiments on two disparate domains confirm the framework’s generality. For a 1000-component building-maintenance task, our method significantly prolongs component survival relative to baseline heuristics and approaches the performance of the oracle policy. On an ETF portfolio-rebalancing problem with draw-down–risk budgets, the same algorithm consistently preserves portfolio viability and outperforms vanilla PPO and the equal-weight baseline. Across both settings, empirical runtimes grow linearly with the number of components, validating the scalability predicted by our complexity analysis. Future work will focus on extending the framework’s capabilities to more dynamic budget allocation schemes and more complicated hierarchical budget constraints.
Appendix A Function Approximation of $\mathbb{E}[T_{\max}]$
We model $\mathbb{E}[T^{i}_{\max}]$ as an exponential function of the budget allocated to component ii. The choice of an exponential function is motivated by its ability to capture the saturation in $\mathbb{E}[T^{i}_{\max}]$ values at higher budget levels, a result of the finite planning horizon $H$. Additionally, the exponential model accounts for non-zero $\mathbb{E}[T^{i}_{\max}]$ even when the budget is zero.
Figure 9: Exponential $\tilde{T}^{i}_{\max}$ curves obtained using non-linear least-squares regression.
To validate the accuracy of this exponential model for $\mathbb{E}[T^{i}_{\max}]$, we conducted non-linear least squares regression on 100 infrastructure components. Figure 9 illustrates the curves obtained through this regression, where $\mathbb{E}[T^{i}_{\max}]$ is modeled as an exponential function. The results indicate that the exponential function provides a strong approximation for $\mathbb{E}[T^{i}_{\max}]$, with an average coefficient of determination $R^{2}_{mean}=0.899$.